A blue car is traveling on the freeway at 55 miles per hou. When first observed, a red car is one half mile behind the blue car. From that point in time, it takes exactly 70 seconds for the red car to overtake and pass the blue car. How fast was the red car traveling?
Answer by mananth(16949)

let speed of red car be x mph
speed of blue car = x+55 mph
catch up distance = 1/2 mile =
catch up time = 70 seconds = 70/3600 hours= 7/360 hours
s= d/t
s= (3/2 ) / (7/360)
s=3/2 * 360/7
s=540/7
x+55 = 540/7
7x+385=540
7x=540-385
7x=155
/7
x=22.14
Speed of red car = 55+22.14 = 77.14 mph